# Foundational theories of hesitant fuzzy sets and hesitant fuzzy information systems and their applications for multi-strength intelligent classifiers

### v3

@@ 2.2 Inclusion relationships of hesitant fuzzy sets
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$ is obtained, where $q=min\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.8 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(8) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for all $x\in U$. By the results of (6) and (7), (8) holds.$\blacksquare$
Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, then $A\subset_{sot}A\cap B$.
(2) If $A\subset_{a}B$, then $A\subset_{sot}A\cap B$.
(3) If $A\subset_{sot}B$, then $A\subset_{sot}A\cap B$.
(4) If $A\subset_{n}B$, then $A\subset_{sot}A\cap B$.
(5) If $A\subset_{sot}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
(6) If $A\subset_{n}B$, then $A\subset_{sot}B\cup C$ for a random $C\in HF(U)$.
Example 2.17.
@@
The following statements hold for $A,B\in HF(U)$,
(1) If $A\subset_{n}B$, $A\cap B\subset_{n}B$.
(2) If $A\subset_{n}B$, $A\subset_{n}A\cup B$.
(3) If $A\subset_{n}B$, $A\cap B\subset_{n}A\cup B$.
Example 2.19.
Let $U=\{x\}$, $A=\frac{\{0.1,0.3,0.5\}}{x}$, $B=\frac{\{0.2,0.4,0.6\}}{x}$.
For three classical sets $\bar{A}$, $\bar{B}$ and $\bar{C}$, if $\bar{A}\sqsubset\bar{B}$, then $\bar{A}\sqcap\bar{C}\sqsubset\bar{B}\sqcap\bar{C}$ and $\bar{A}\sqcup\bar{C}\sqsubset\bar{B}\sqcup\bar{C}$. These rules in the cases of three hesitant fuzzy sets $A$, $B$ and $C$ are investigated, shown as Proposition 2.20.
@@
$v_{b}=B^{-}(x)\leqslant(B\cup C)^{-}(x)$, then $(B\cup C)(x)$ is a subsequence of $W^{2}$. Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
@@
$o_{a}=A^{-}(x)\leqslant(A\cup C)^{-}(x)$, then $(A\cup C)(x)$ is a subsequence of $W^{1}$. Let $q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=\min\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
@@
(13) By Proposition 2.9 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for $x\in U$. By the results of (11) and (12), then (13) holds.$\blacksquare$
(11) and (12) of Proposition 2.20 can be combined as (1) of Proposition ${}^{\prime}$. (13) of Proposition 2.20 can be written into (2) of Proposition ${}^{\prime}$.
Proposition ${}^{\prime}$ The following statements hold for $A,B,C\in HF(U)$,
@@
Proof (1) For all $x\in U$, $A^{+}(x)\leqslant B^{+}(x)$ and $A^{+}(x)\leqslant C^{+}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{+}(x),C^{+}(x)\}$.
(2) For all $x\in U$, $A^{-}(x)\leqslant B^{-}(x)$ and $A^{-}(x)\leqslant C^{-}(x)$ if and only if $A^{-}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$. To combine the result of (1), then (2) holds.
@@
(4) For all $x\in U$, $A^{+}(x)\leqslant B^{-}(x)$ and $A^{+}(x)\leqslant C^{-}(x)$ if and only if $A^{+}(x)\leqslant\min\{B^{-}(x),C^{-}(x)\}$$\blacksquare$
@@
Definition 2.23.
TV For $x\in U$ and $H\in HF(U)$, the complement of $H$ is denoted as $H^{c}$, where
$H^{c}(x)=\sqcup_{\gamma\in H(x)}\{1-\gamma\}$.
Theorem 2.24.
TV ; XX The following statements hold for $A,B,C\in HF(U)$,
(1) $(A^{c})^{c}=A$.
(2) $(A\cap B)^{c}=A^{c}\cup B^{c}$.
(3) $(A\cup B)^{c}=A^{c}\cap B^{c}$.
(4) $A\cap B=B\cap A$, $A\cup B=B\cup A$.
(5) $(A\cap B)\cap C=A\cap(B\cap C)$, $(A\cup B)\cup C=A\cup(B\cup C)$.
@@
Let $U=\{x,y\}$, $A=\frac{0.4,0.4}{x}+\frac{0.2,0.25}{y}$, $B=\frac{0.1,0.1,0.41}{x}+\frac{0.1,0.2,0.3}{y}$.
@@
2.3 Equivalence relationships of hesitant fuzzy sets
@@
For the proofs of (4)-(7), we give a simple example as follows. By the results of fuzzy sets ZADE , $(A^{\prime}(x)\cup B^{\prime}(x))\cap A^{\prime}(x)=A^{\prime}(x)$ and $(A^{\prime\prime}(x)\cup B^{\prime\prime}(x))\cap A^{\prime\prime}(x)=A^{\prime\prime}(x)$, i.e., $(A^{+}(x)\cup B^{+}(x))\cap A^{+}(x)=A^{+}(x)$ and $(A^{-}(x)\cup B^{-}(x))\cap A^{-}(x)=A^{-}(x)$, then $(A\cup B)\cap A=_{a}A$.
(4)-(7) are easy to be proved by the results of fuzzy sets in ZADE $\blacksquare$
@@
Let $U=\{x\}$, $A=\frac{0.1,0.2,0.3}{x}$, $B=\frac{0.3,0.4,0.5}{x}$, $C=\frac{0.3,0.45,0.5}{x}$.
@@
For the $\subset_{p}$ and $=_{p}$ relationships, the upper bound with respect to an object $x\in U$ and a hesitant fuzzy set over $U$ is the unique thing that needs to be investigated. Hence, many inclusion relationships and equivalence relationships of fuzzy sets also hold on hesitant fuzzy sets as the forms of $\subset_{p}$ and $=_{p}$ relationships. However, $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}$, then the $\subset_{p}$ and $=_{p}$ relationships are not consistent with the inclusion relationships and equivalence relationships of fuzzy sets.
Section 2.2 and 2.3 reveal that some common inclusion relationships and equivalence relationships do not hold on hesitant fuzzy sets but hold on classical sets.
@@ 3 Foundations of families of hesitant fuzzy sets
If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, let $V=(\bigcap\{H:H\in\mathscr{H}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}\})(x)$, $min((\bigcup\{H:H\in\mathscr{H}\})(x)-V)>H_{inf}^{+}(x)>H_{sup}^{-}(x)>max((\bigcap\{H:H\in\mathscr{H}\})(x)-V)$, by Remark 2.8 (4), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$$\blacksquare$
@@
Let $V=(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. Since $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $sup_{1}^{-}\leqslant sup_{2}^{-}$. If $h\in(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. It means that $max((\bigcup\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$. On the other hand, $sup_{2}^{-}<min((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.8 (4), then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.

### v4

@@ 2.3 Some propositions about the intersection and union of two hesitant fuzzy sets
Based on the illustrations of sub-case $B^{-}(x)<C^{-}(x)$, $A(x)(q)\preccurlyeq(B\cup C)(x)(q)$ is obtained, where $q=inf\{|A(x)|,|(B\cup C)(x)|\}$, by Remark 2.11 (3), one of $A(x)\subset_{s}(B\cup C)(x)$ and $A(x)\subset_{t}(B\cup C)(x)$ holds.
(5) By Proposition 2.12 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for all $x\in U$. By the results of (3) and (4) of Proposition 2.22, (5) holds.$\blacksquare$
(3) and (4) of Proposition 2.22 can be combined as (1) of Proposition 2.22. (5) of Proposition 2.22 can be written into (2) of Proposition 2.22.
Proposition 2.22 The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{sot}B$, then $A\subset_{sot}B\cup C$.
(2) If $A\subset_{n}B$, then $A\subset_{sot}B\cup C$.
@@
The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$, $A\cap C\subset_{p}B\cap C$.
(2) If $A\subset_{a}B$, $A\cap C\subset_{a}B\cap C$.
(3) If $A\subset_{s}B$, $A\cap C\subset_{a}B\cap C$.
(4) If $A\subset_{t}B$, $A\cap C\subset_{p}B\cap C$.
(5) If $A\subset_{n}B$, $A\cap C\subset_{a}B\cap C$.
Proof (1)-(5) are obvious.
@@
$v_{b}=B^{-}(x)\leqslant(B\cup C)^{-}(x)$, then $(B\cup C)(x)$ is a subsequence of $W^{2}$. Let $q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
@@
$o_{a}=A^{-}(x)\leqslant(A\cup C)^{-}(x)$, then $(A\cup C)(x)$ is a subsequence of $W^{1}$. Let $q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, then $q\leqslant|W^{1}|=|W^{2}|$. $w^{1}_{j}\leqslant w^{2}_{j}$ for $1\leqslant j\leqslant q=inf\{|(A\cup C)(x)|,|(B\cup C)(x)|\}$, where $w^{1}_{j}\in W^{1}$ and $w^{2}_{j}\in W^{2}$.
@@
(8) By Proposition 2.12 (9), $A\subset_{n}B$ implies $A(x)\subset_{s}B(x)$ or $A(x)\subset_{t}B(x)$ for $x\in U$. By the results of (6) and (7) of Proposition 2.24, then (8) holds.$\blacksquare$
(6) and (7) of Proposition 2.24 can be combined as (1) of Proposition 2.24. (8) of Proposition 2.24 can be written into (2) of Proposition 2.24.
Proposition 2.24 The following statements hold for $A,B,C\in HF(U)$,
@@
Proof (1) For all $x\in U$, $A^{+}(x)\leqslant B^{+}(x)$ and $A^{+}(x)\leqslant C^{+}(x)$ if and only if $A^{+}(x)\leqslant inf\{B^{+}(x),C^{+}(x)\}$.
(2) For all $x\in U$, $A^{-}(x)\leqslant B^{-}(x)$ and $A^{-}(x)\leqslant C^{-}(x)$ if and only if $A^{-}(x)\leqslant inf\{B^{-}(x),C^{-}(x)\}$. To combine the result of (1), then (2) holds.
@@
(4) For all $x\in U$, $A^{+}(x)\leqslant B^{-}(x)$ and $A^{+}(x)\leqslant C^{-}(x)$ if and only if $A^{+}(x)\leqslant inf\{B^{-}(x),C^{-}(x)\}$$\blacksquare$
@@
2.5 The proposition of the complement of hesitant fuzzy set and the transitivity of the inclusion relationships in hesitant fuzzy sets
@@
Let $U=\{x,y\}$, $A=\frac{\{0.4,0.4\}}{x}+\frac{\{0.2,0.25\}}{y}$, $B=\frac{\{0.1,0.1,0.41\}}{x}+\frac{\{0.1,0.2,0.3\}}{y}$.
@@
Proposition 2.29.
The following statements hold for $A,B,C\in HF(U)$,
(1) If $A\subset_{p}B$ and $B\subset_{p}C$, then $A\subset_{p}C$.
(2) If $A\subset_{a}B$ and $B\subset_{a}C$, then $A\subset_{a}C$.
(3) If $A\subset_{m}B$ and $B\subset_{m}C$, then $A\subset_{m}C$.
(4) If $A\subset_{s}B$ and $B\subset_{s}C$, then $A\subset_{s}C$.
(5) If $A\subset_{t}B$ and $B\subset_{t}C$, then $A\subset_{t}C$.
(6) If $A\subset_{n}B$ and $B\subset_{n}C$, then $A\subset_{n}C$.
Proof (1) and (2) are obvious.
2.6 The situation of equality in hesitant fuzzy sets
@@
For the proofs of (4)-(7), we give a simple example as follows. By the results of fuzzy sets [6], $(A^{\prime}(x)\cup B^{\prime}(x))\cap A^{\prime}(x)=A^{\prime}(x)$ and $(A^{\prime\prime}(x)\cup B^{\prime\prime}(x))\cap A^{\prime\prime}(x)=A^{\prime\prime}(x)$, i.e., $(A^{+}(x)\cup B^{+}(x))\cap A^{+}(x)=A^{+}(x)$ and $(A^{-}(x)\cup B^{-}(x))\cap A^{-}(x)=A^{-}(x)$, then $(A\cup B)\cap A=_{a}A$.
(4)-(7) are easy to be proved by the results of fuzzy sets in [6]$\blacksquare$
@@
Let $U=\{x\}$, $A=\frac{\{0.1,0.2,0.3\}}{x}$, $B=\frac{\{0.3,0.4,0.5\}}{x}$, $C=\frac{\{0.3,0.45,0.5\}}{x}$.
@@
When considering the $\subset_{p}$ and $=_{p}$ relationships, the primary focus lies on determining the upper bound for an object $x\in U$ and a hesitant fuzzy set defined over $U$. Consequently, several inclusion and equal situations that are applicable to fuzzy sets can also be extended to hesitant fuzzy sets in the form of $\subset_{p}$ and $=_{p}$ relationships. Nevertheless, the absence of the implication $A\subset_{p}B\nRightarrow B^{c}\subset_{p}A^{c}$ indicates that the $\subset_{p}$ and $=_{p}$ relationships do not align with the inclusion relationships and equal situations observed in fuzzy sets.
Sections 2.2 and 2.3 reveal that certain commonly observed inclusion and equal situations do not apply to hesitant fuzzy sets but are valid for classical sets.
@@ 3 Foundations of families of hesitant fuzzy sets
If $H_{sup}^{-}(x)<H_{inf}^{+}(x)$, let $V=(\bigcap\{H:H\in\mathscr{H}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}\})(x)$, $inf((\bigcup\{H:H\in\mathscr{H}\})(x)-V)>H_{inf}^{+}(x)>H_{sup}^{-}(x)>sup((\bigcap\{H:H\in\mathscr{H}\})(x)-V)$, by Remark 2.11 (4), ($(\bigcap\{H:H\in\mathscr{H}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}\})(x)$$\blacksquare$
@@
Let $V=(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\sqcap(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. Since $\mathscr{H}_{1}\sqsubset\mathscr{H}_{2}$, then $sup_{1}^{-}\leqslant sup_{2}^{-}$. If $h\in(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)$ and $sup_{2}^{-}\leqslant h$, then $h\in(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$. It means that $sup((\bigcup\{H:H\in\mathscr{H}_{1}\})(x)-V)<sup_{2}^{-}$. On the other hand, $sup_{2}^{-}<inf((\bigcup\{H:H\in\mathscr{H}_{2}\})(x)-V)$, by Remark 2.11 (4), then ($(\bigcup\{H:H\in\mathscr{H}_{1}\})(x)\subset_{sot}(\bigcup\{H:H\in\mathscr{H}_{2}\})(x)$.